ISRAEL JOURNAL OF MATHEMATICS

TABLE OF CONTENTS, VOLUME 198, 2013

Brooks, S. Partially localized quasimodes in large subspaces  393

Cabello Sánchez, F. and J. M. F. Castillo Stability constants and the homology of quasi-Banach spaces

347

Casolo, C. and O. Puglisi Nil-automorphisms of groups with residual properties

91

Castillo, J. M. F. and Y. Moreno On the bounded approximation property in Banach spaces

243

Castillo, J. M. F. See F. Cabello Sánchez Chen, G. and S. Ma Ground state periodic solutions of second order

Hamiltonian systems without spectrum 0

111

Cheng, C. Multiplicity one of regular Serre weights  419

Conley, C. T., A. S. Kechris and B. D. Miller Stationary probability measures and topological realizations

333

Crespo, T., Z. Hajto and E. Sowa Picard–Vessiot theory for real fields

75

Crestani, E. and A. Lucchini The generating graph of finite soluble groups

63

Dolfi, S., G. Navarro and P. H. Tiep Finite groups whose same degree characters are Galois conjugate

283

Essert, J. Homological stability for classical groups  169

Felmer, P. and E. Topp Convergence results for a class of nonlinear fractional heat equations

1

Hajto, Z. See T. Crespo Huang, M., Y. Li, M. Vuorinen and X. Wang On quasimöbius

maps in real Banach spaces

467

Huang, Z. Y. Gorenstein injective and strongly cotorsion modules  215

Im, B.-H. and M. Larsen Some applications of the Hales–Jewett theorem to field arithmetic

35

Kechris, A. S. See C. T. Conley Larsen, M. See B.-H. Im Lei, A., D. Loeffler and S. L. Zerbes Critical slope p-adic L-

functions of CM modular forms

261

Li, Y. See M. Huang Loeffler, D. See A. Lei Lucchini, A. See E. Crestani Ma, S. See G. Chen Melleray, J. and T. Tsankov Generic representations of abelian

groups and extreme amenability

129

Menezes, N. E., M. Quick and C. M. Roney-Dougal The probability of generating a finite simple group

371

Miller, B. D. See C. T. Conley Moreno, Y. See J. M. F. Castillo Navarro, G. See S. Dolfi Pinchasi, R. A solution to a problem of Grünbaum and Motzkin and

of Erdős and Purdy about bichromatic configurations of points in the plane

205

Puglisi, O. See C. Casolo Quick, M. See N. E. Menezes Roney-Dougal, C. M. See N. E. Menezes Sarkar, R. P. Chaotic dynamics of the heat semigroup on the

Damek–Ricci spaces

487

Sowa, E. See T. Crespo

Stix, J. Birational p-adic Galois sections in higher dimensions  49

Tân, N. D. Embedding problems with local conditions and the admissibility of finite groups

229

Tiep, P. H. See S. Dolfi Topp, E. See P. Felmer Tsankov, T. See J. Melleray Vuorinen, M. See M. Huang Wang, X. See M. Huang Zerbes, S. L. See A. Lei